OpenStudy (anonymous):

Please help!! Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

terenzreignz (terenzreignz):

Verify or prove? :) Anyway, use these identities, which are actually practically the same identity :) \[\large \cos(x+y)=\cos \ x\cos \ y - \sin \ x\sin\ y \]\[\large \cos(x-y)=\cos \ x\cos \ y+\sin \ x \sin\ y\]

OpenStudy (anonymous):

multuply the first equation by -1, then add two equation, you will find the equation that you are looking for..